Az április 4-i dolgozat eredményei.

feladatok  1.  2.  3.  4a. 4b.  5  6.

diákok:

Andrékó Judit  −  −  −  −  +  −
Bácskai Marianna  +  +  +  +  −  ±  ±
Balázs Erika  +  +  ±  +  ±  +  ±
Bálint Melinda  +  +  +  +  ±  +  −
Balogh Péter  −  −  −  +  −
Benczik Attila  −  −  −  −  −  −
Boros Ildikó  +  +  +  +  −  +  ±
Czibere Marianna  −  −  −  +  −  ±  ±
Császár Katalin  −  ±  +  +  ±  +  ±
Dienes Zoltán  −  −  −  +  −  −  −
Éliás Anikó  +  +  +  + −  +  ±

Fórizs Fruzsina

Föczén Beáta  +  +  ±  +  ±  +  +
Gőgh Timea  +  +  −  −  −  +  ±

Gombkötő Angéla

Holló Anikó  +  +  +  −  −  ±
Huszty Árpád  ±  −  −  −  −  −
Iván Éva  −  −  −  −  −  −
Kelemen Imre  +  +  +  +  −  +  ±
Kerékgyártó Zita  ±  −  −  +  −  +
Kocsis Zsuzsanna  ±  +  ±  ±  −
Koncz Nikoletta  +  ±  +  +  ±  ±  +
Lovas Rezső László  +  +  +  +  +  +  +
Lukács Erika  +  +  +  −  ±  −
Orosz Erzsébet  −  +  −  −  −  +  −
Pelei Csaba  +  ±  +

Pintér Györgyi

Szabó Erika  +  +  +  +  −  +  ±
Szilágyi Renáta Judit  −  −  ±  −  −  ±  −
Sztankó Paticia  −  ±  −  +  −  ±
Szűcs Mónika  +  −  +  +  −  ±  ±
Turák Olivér  −  +  ±  +  −  −  ±
Villás Rita  +  +  ±  + ±  +  +